### v1

# Optimization for expectation value estimation with shallow quantum circuits

Abstract
Estimating linear properties of quantum states, such as fidelities, molecular energies, and correlation functions, is a fundamental task in quantum information science. The classical shadow has emerged as a prevalent tool due to its efficiency in estimating many independent observables simultaneously. However, it does not utilize the information of the target observable and the constraints of quantum devices, making it inefficient in many practical scenarios where the focus is on estimating a select few observables. To address this inefficiency, we propose a framework that optimizes sample complexity for estimating the expectation value of any observable using a shallow parameterized quantum circuit. Within this framework, we introduce a greedy algorithm that decomposes the target observable into a linear combination of multiple observables, each of which can be diagonalized with the shallow circuit. Using this decomposition, we then apply an importance sampling algorithm to estimate the expectation value of the target observable. We numerically demonstrate the performance of our algorithm by estimating the ground energy of a sparse Hamiltonian and the inner product of two pure states, highlighting the advantages compared to some conventional methods. Additionally, we derive the fundamental lower bound for the sample complexity required to estimate a target observable using a given shallow quantum circuit, thereby enhancing our understanding of the capabilities of shallow circuits in quantum learning tasks.
I Introduction
Learning the properties of an unknown quantum state is crucial in the fields of many-body physics, quantum chemistry, and quantum information science [1, 2, 3, 4, 5, 6]. Among these properties, the linear properties $\text{Tr}\left(\rho H\right)$, where $H$ represents an observable, are of special importance since quantum mechanics is inherently linear. According to the fundamental principles of quantum mechanics, any information that can be directly extracted from a quantum system is linear, and complex properties must be calculated based on these linear properties. Therefore, various protocols have been developed to facilitate the estimation of linear properties, ranging from conventional tomography-based protocols [7, 8, 9] and Pauli decomposition methods [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21] to quantum memory-assisted protocols [22, 23]. Moreover, understanding the fundamental challenges in estimating linear properties can deepen our comprehension of quantum learning problems [24, 23, 25, 26].
Among all the estimation protocols, randomized measurement [27] has gradually become a prevailing toolbox due to its experimental accessibility [28, 29, 30, 31] and efficiency in estimating certain state properties [32, 33, 34, 35, 36, 37, 38]. Notably, Huang et al. introduced the classical shadow (CS) protocol [39], which aims to provide a classical description of an unknown state $\rho$ and use it to estimate arbitrary state properties. Conducted in a “measure first, ask questions later” manner [27], the CS protocol does not require the information of observables during the experimental stage and can efficiently estimate a large number of independent observables simultaneously. However, aside from local observables, the original CS protocol relies on random Clifford unitaries to achieve low sample complexity, necessitating a deep quantum circuit. Furthermore, since the quantum experiment does not utilize the information of target observables, the CS protocol is inefficient when the target is to estimate only a select few observables.
Since the CS protocol was proposed, considerable effort has been devoted to improving its performance in practical scenarios and addressing its limitations. To make the CS protocol more suitable for noisy intermediate-scale quantum devices, a random shallow circuit-based CS protocol was developed [40, 41, 42]. With fewer quantum gates, the shallow CS protocol achieves comparable, and sometimes better, performance than the random Clifford CS protocol for certain non-local observables [43]. However, the shallow CS protocol remains inefficient for estimating a small number of observables since the experimental stage is still independent of the target observables. Various methodologies have been developed to improve the efficiency of the CS protocol by leveraging information about target observables [10, 11, 12, 13, 14, 15, 16, 17, 19, 20, 18, 21]. Given the Pauli decomposition $H=\sum_{i=1}^{m}\alpha_{i}Q_{i}$, these methods aim to efficiently estimate the expectation values for all $\{Q_{i}\}_{i=1}^{m}$. This is achieved by performing Clifford circuits sampled from a specific set $\mathcal{S}$ with certain probabilities, followed by measurement in the computational basis and classical post-processing. The gate set $\mathcal{S}$ and these probabilities are carefully designed and vary depending on the specific algorithms. While these methods improve performance compared to the original CS protocol by utilizing information about target observables, they do not consider the practical constraints of quantum circuits. Firstly, implementing some target Clifford gates may require deep circuits [11, 21]. Additionally, since the observable decomposition is restricted to Pauli observables, these methods may still require high sample complexity even with deep circuits.
In this work, we propose a framework for estimating linear properties that unifies the logic of the methodologies mentioned earlier. Specifically, given an unknown quantum state and a shallow variational circuit determined by experimental settings, our framework aims to use the information of the target observable to optimize the usage of the shallow circuit and achieve low sample complexity. Our approach is inspired by the observation that, given an observable $H$, if we can perform any unitary operation on the quantum state $\rho$, $\mathcal{O}(\left\|H\right\|_{2}\epsilon^{-2})$ copies of $\rho$ are sufficient to approximate $\text{Tr}\left(\rho H\right)$ with a precision of $1-\epsilon$ and high success probability, where $\|\cdot\|_{2}$ denotes the spectral norm. This requirement arises from the eigen-decomposition of $H$, given by $H=U^{\dagger}\Lambda U$. Thus, we can rotate the quantum state using the unitary $U$ and measure it in the computational basis to estimate $\text{Tr}\left(\rho H\right)$. In practical scenarios, where we are limited to $L$-depth parameterized unitaries $U_{L}(\theta)$, we can decompose the target observable as $H\approx\sum_{k=1}^{K}U_{L}(\theta^{(k)})^{\dagger}\Lambda_{k}U_{L}(\theta^{(k)})$, with $\Lambda_{k}$ being real diagonal matrices. Each term in the summation is an observable that can be diagonalized with the parameterized circuit to estimate $\text{Tr}\left(U_{L}(\theta^{(k)})^{\dagger}\Lambda_{k}U_{L}(\theta^{(k)})\rho\right)$. By optimizing this decomposition, we can achieve low sample complexity.
Based on these observations, our framework can be divided into classical and quantum phases. In the classical phase, we use an algorithm to decompose the target observable within the constraints of the available quantum circuit, including its depth and structure. In the quantum phase, we design an efficient method to estimate the expectation value of the target observable based on the given decomposition. Many existing Pauli decomposition protocols, such as qubit-commuting measurement approaches [10, 13, 15, 18] and commuting measurement approaches [11, 21], fall within the scope of our framework. Within this framework, we adopt a greedy algorithm to numerically find the decomposition and employ an importance sampling algorithm to estimate $\text{Tr}\left(\rho H\right)$, as illustrated in Fig. 1. The required number of copies of $\rho$ is $\mathcal{O}\left(\frac{\|\bm{\Lambda}\|_{1}^{2}\log(1/\delta)}{\epsilon^{2}}\right)$ with precision $1-\epsilon$ and success probability $1-\delta$, where $\bm{\Lambda}$ is a $K$-dimensional vector with the $k$-th entry being $\bm{\Lambda}(k)=\|\Lambda_{k}\|_{2}$, and $\left\|\bm{\Lambda}\right\|_{1}$ denotes the sum of the absolute values of all the entries in $\bm{\Lambda}$.
To verify the convergence of our observable decomposition algorithm, we numerically evaluate the distance between the approximate observable $\sum_{k=1}^{K}U_{L}(\theta^{(k)})^{\dagger}\Lambda_{k}U_{L}(\theta^{(k)})$ and the target observable across various circuit depths $L$ and the number of decomposition terms $K$. The numerical results indicate that for a fixed number of decomposition terms, this distance decreases as the depth of the ansatz circuits increases, which aligns with our intuition. Moreover, the Frobenius distance decays exponentially as the number of decomposition terms increases when fixing the circuit depth. We further demonstrate the practical efficiency of our algorithms by applying them to estimate the ground energy of a sparse Hamiltonian and the inner product of two pure quantum states, comparing the results with conventional methods.
In addition to providing a practical measurement framework, we also aim to address the fundamental question: What is the limitation of a shallow quantum circuit in estimating an observable? For a given variational quantum circuit $U_{L}(\theta)$, we establish a lower bound on the number of copies of the quantum state required to accurately estimate $\text{Tr}\left(\rho H\right)$. This bound is given by $\Omega\left(\frac{\text{Tr}\left(H_{0}^{2}\right)^{2}}{\epsilon^{2}\delta(H_{0})4^{n}}\right)$, where $H_{0}$ represents the traceless part of $H$, and $\delta(H_{0})$ is the square of the maximum expectation value of $H_{0}$ with respect to the quantum states prepared by applying $U_{L}(\theta)^{\dagger}$ to a product state. We also analyze the tightness of this lower bound in specific cases.
II Expectation value estimation framework
Our framework can be roughly divided into two phases. The classical phase relies on a specific decomposition algorithm to decompose the target observable into several terms. In the quantum phase, we estimate the expectation value based on the decomposition of the target observable and some probability distribution. We summarize the framework using Algorithm 1, in which $\left\|\cdot\right\|_{2}$ denotes the spectral norm, $\left\|A\right\|_{2}=\max_{j}\sigma_{j}$ with $\sigma_{j}$ being the singular value of matrix $A$, $\bm{\Lambda}$ represents a $K$-dimensional vector with the $k$-th entry being $\bm{\Lambda}(k)=\left\|\Lambda_{k}\right\|_{2}$, and $\left\|\bm{\Lambda}\right\|_{1}$ denotes the sum of the absolute values of all the entries in $\bm{\Lambda}$.
input : Quantum state $\rho$, observable $H$, error $\epsilon_{1}$, the number of samplings $T$, and probabilities $\left\{p_{k}\right\}_{k=1}^{K}$.
output : Estimator of $\text{Tr}\left(\rho H\right)$.
1 Generate unitaries $\left\{U_{k}\right\}_{k=1}^{K}$ and real diagonal matrices $\left\{\Lambda_{k}\right\}_{k=1}^{K}$ such that $\left\|H-\sum_{k=1}^{K}U_{k}^{\dagger}\Lambda_{k}U_{k}\right\|_{2}\leq\epsilon_{1}$;
2 for $t\leftarrow 1$ to $T$ do
3      Sample an index $k$ with probability $p_{k}$;
4      Perform $U_{k}$ to the prepared state $\rho$, then measure in computational basis to obtain the $n$-bit measurement result $b_{t}$;
5      Let the estimator $v_{t}=\frac{\Lambda_{k}(b_{t})}{p_{k}}$;
6
7      end for
return $v:=\operatorname{\mathbf{MedianOfMeans}}(v_{1},\cdots,v_{t})$
Algorithm 1 Framework for Estimating Linear Properties of Quantum States.
The central process of the classical stage is generating the decomposition $\{U_{k},\Lambda_{k}\}_{k}$ of the target observable. In fact, many Pauli decomposition-based quantum measurement protocols offer methodologies for generating such decompositions. Below, we will illustrate two of these methods: one based on qubit-commuting Pauli decomposition and the other on commuting Pauli decomposition.
  • (1)
    Qubit-commuting decomposition. We call Pauli operators $P$ and $Q$ qubit-commuting if, for any $l$-th qubit, $[P_{l},Q_{l}]=0$. Since the $4^{n}$ Pauli operators form a basis for the space of $2^{n}\times 2^{n}$ Hermitian matrices, any observable $H$ can be expressed as $H=\sum_{i=1}^{m}\alpha_{i}Q_{i}$, where $\{\alpha_{i}\}_{i}$ are real coefficients and $\{Q_{i}\}_{i}$ are Pauli operators. Since qubit-commuting Pauli observables can be estimated simultaneously, we can group the Pauli operators $Q_{1},\ldots,Q_{m}$ into $K$ qubit-commuting (overlapping) sets $e_{1},e_{2},\ldots,e_{K}$, where each set contains Pauli operators that commute with each other. Various grouping methods have been proposed, such as the largest degree first algorithm [12] and the overlapping grouping algorithm [18]. Given a grouping algorithm, we can express the target observable as $H=\sum_{k=1}^{K}U_{k}^{\dagger}\Lambda_{k}U_{k}$, where each $U_{k}$ is the tensor product of local Clifford gates associated with the eigenstates of Pauli operators in the set $e_{k}$, and the diagonal matrix $\Lambda_{k}$ contains information about the eigenvalues of the Pauli operators in $e_{k}$. Further details on the explicit representations of $U_{k}$ and $\Lambda_{k}$ are provided in the Appendix B.
  • (2)
    Commuting decomposition. The qubit-commuting sets can be relaxed to the fully commuting sets $\left\{g_{k}\right\}_{k=1}^{K}$, where any two Pauli operators in $g_{k}$ commute with each other. Consequently, the target observable can be expressed as $H=\sum_{k=1}^{K}H_{k}$, with $H_{k}=\sum_{Q_{j}\in g_{k}}\alpha_{j}Q_{j}$. Similarly, there exists a set of Clifford gates $\{U_{k}\}_{k}$, which may require a polynomial-size quantum circuit to implement, and diagonal matrices $\{\Lambda_{k}\}_{k}$ such that $H_{k}=U_{k}^{\dagger}\Lambda_{k}U_{k}$ [11] for all $k$. Details on how to generate $\{U_{k},\Lambda_{k}\}_{k}$ are provided in the Appendix D.
The central objective of the quantum stage of our framework is to determine a probability distribution $\{p_{k}\}_{k}$ used to sample the estimated observables. In Pauli decomposition-based measurement protocols, this probability distribution corresponds to the one used for sampling different Pauli groups. Wu et al. [18] demonstrated that all existing qubit-commuting protocols can be unified through two central processes: Pauli grouping and assigning each group a probability, which fits precisely within our framework. Similarly, we show in the Appendix D that existing commuting measurement schemes [11, 21] also fit within this framework. It is important to note that Pauli decomposition-based protocols are limited to Clifford unitary evolution, which cannot achieve the optimal performance for a general observable even with arbitrary Clifford unitaries. Therefore, our goal is to design new algorithms that are compatible with general quantum circuits.
III Expectation value estimation with shallow circuit
Figure 1: Overview of the expectation value estimation with shallow circuit. We utilize a classical greedy algorithm to generate a decomposition of $H\approx\sum_{k=1}^{K}U_{L}(\theta^{(k)})^{\dagger}\Lambda_{k}U_{L}(\theta^{(k)})$, followed by iteratively repeating $T$ rounds: sampling $k$ with probability $p_{k}\propto\left\|\Lambda_{k}\right\|_{2}$, performing $U_{L}(\theta^{(k)})$ on the prepared quantum state $\rho$, measuring in computational basis with result $b$, collecting $\Lambda_{k}(b)/p_{k}$, to generate estimation $v$, where $\Lambda_{k}(b)$ denotes the $(b,b)$-th element of the diagonal matrix $\Lambda_{k}$ and $b$ ranges from 00 to $12^{n}-$ 1.
Based on the introduced framework, we design specific algorithms to generate the decomposition of the target observable and estimate the expectation value, as illustrated in Fig. 1. We consider a practical scenario, where the depth and structure of the quantum circuit are given according to the capabilities of the available quantum devices and can go beyond Clifford circuits. Without loss of generality, we hereafter take a typical circuit structure: an $L$-depth parameterized quantum circuit $U_{L}(\theta)$ with the brick-like arrangement of 2-qubit parametrized gates, to demonstrate our algorithms. The 0-layer of the ansatz circuit comprises a layer of arbitrary single-qubit gates. This parameterized circuit is tomography-complete for any $L\geq 0$ as it can estimate all Pauli observables, even when $L=0$. Our conclusions can be easily generalized to other circuit structures.
We propose a greedy algorithm to decompose the target observable $H$ using the $L$-depth parameterized circuit $U_{L}(\theta)$. In each step, we aim to find an observable, $U_{L}(\theta)^{\dagger}\Lambda U_{L}(\theta)$, that approximates $H$. This involves finding a unitary $U_{L}(\theta)$ and a real diagonal matrix $\Lambda$ that minimize the distance between $U_{L}(\theta)HU_{L}(\theta)^{\dagger}$ and $\Lambda$. As shown in Appendix A, when using the Frobenius norm to quantify the distance, the optimal choice of $\Lambda$ is $\Lambda=\mathrm{diag}[U_{L}(\theta)HU_{L}(\theta)^{\dagger}]$, where $\mathrm{diag}(X)$ denotes a diagonal matrix whose elements are the diagonal elements of $X$. As a comparison, the distance based on the 2-norm brings us an additional optimization step to determine the optimal $\Lambda$. This iterative process updates the Hamiltonian $H$ as outlined in Algorithm 2.
input : Observable $H$, $L$-depth parameterized quantum circuit $U_{L}(\theta)$, accuracy $\epsilon$.
output : Set $\left\{\theta^{(k)},\Lambda_{k}\right\}_{k}$ such that $\left\|H-\sum_{k}U_{L}(\theta_{k})^{\dagger}\Lambda_{k}U_{L}(\theta_{k})\right\|_{2}\leq\epsilon$.
1 $H^{(0)}=H$;
2 $k=0$;
3 while $\left\|H^{(k)}\right\|_{2}\geq\epsilon$ do
4      $\theta^{(k)}=\arg\min_{\theta}\left\|U_{L}\left(\theta\right)H^{(k)}U_{L}^{\dagger}\left(\theta\right)-\mathrm{diag}\left[U_{L}\left(\theta\right)H^{(k)}U_{L}^{\dagger}\left(\theta\right)\right]\right\|_{F}$;
5
6     $\Lambda_{k}=\mathrm{diag}\left[U_{L}\left(\theta^{(k)}\right)H^{(k)}U_{L}^{\dagger}\left(\theta^{(k)}\right)\right]$;
7      $H^{(k+1)}=H^{(k)}-U_{L}^{\dagger}\left(\theta^{(k)}\right)\Lambda_{k}U_{L}\left(\theta^{(k)}\right)$;
8      $k=k+1$;
9
10      end while
return $\left\{\theta^{(k)},\Lambda_{k}\right\}$
Algorithm 2 Greedy Projected Decomposition Algorithm.
Using the decomposition generated by Algorithm 2 and setting $p_{k}\propto\left\|\Lambda_{k}\right\|_{2}$, we claim that Algorithm 1 can produce an estimator of $\text{Tr}\left(\rho H\right)$ with bounded error.
Theorem 1.
Given that $U_{L}(\theta)$ is tomography-complete. For any Hermitian matrix $H$ and integer $L\geq 0,\epsilon_{1},\epsilon_{2}\in(0,1)$, and $\epsilon_{1}+\epsilon_{2}\leq 1$, Algorithm 2 can terminate with a finite set $\{\theta^{(k)},\Lambda_{k}\}_{k=1}^{K}$ satisfying
$\left\|H-\sum_{k=1}^{K}U_{L}\left(\theta^{(k)}\right)^{\dagger}\Lambda_{k}U_{L}\left(\theta^{(k)}\right)\right\|_{2}\leq\epsilon_{1}$ (1)
for some finite integer $K$. Given this decomposition and setting $p_{k}\propto\left\|\Lambda_{k}\right\|_{2}$, Algorithm 1 requires $T=\mathcal{O}\left(\frac{\left\|\bm{\Lambda}\right\|_{1}^{2}\log(1/\delta)}{\epsilon_{2}^{2}}\right)$ copies of $\rho$ to estimate $\text{Tr}\left(\rho H\right)$ to accuracy $\epsilon=\epsilon_{1}+\epsilon_{2}$ with success probability of at least $1-\delta$. Here, $\bm{\Lambda}$ is a $K$-dimensional vector with the $k$-th entry being $\left\|\Lambda_{k}\right\|_{2}$ by choosing $p_{k}\propto\left\|\Lambda_{k}\right\|_{2}$.
Here, the total error $\epsilon$ depends on both the approximation error of $H$ and the estimation error of $\text{Tr}\left(\rho\widehat{H}\right)$, where $\widehat{H}=\sum_{k=1}^{K}U_{L}\left(\theta^{(k)}\right)^{\dagger}\Lambda_{k}U_{L}\left(\theta^{(k)}\right)$. To minimize the resource consumption, we aim to minimize the value of $K$ while keeping a small approximation error. In the following numerical results, we will demonstrate that our algorithms, using a shallow parameterized circuit, require significantly fewer values of $K$ to decompose a sparse Hamiltonian with some error, compared to the number of terms with the Pauli decomposition approach.
IV Numerical results
We evaluate the performance of Algorithm 1 in combination with Algorithm 2 and importance sampling using two examples. In the first example, we estimate the expectation value of randomly generated sparse Hamiltonians. The second example involves estimating the inner product of a Slater determinant |$\ket{\phi_{\tau}}$ with a given pure state |$\ket{\psi}$, denoted as $\braket{\phi_{\tau}}{\psi}$. In our numerical experiments, each two-qubit parameterized gate is an iSWAP gate followed by the tensor product of two arbitrary single-qubit gates. In each example, we first demonstrate the convergence of Algorithm 2 as the number of decomposition terms of $H$ and the depth $L$ of the ansatz $U_{L}(\theta)$ increase. We then use the importance sampling probability distribution to sample the decomposed observables and compare the sample complexity of our algorithms with several existing methods [13, 18, 20]. As a subroutine, Algorithm 2 requires optimizing $\left\|U_{L}\left(\theta\right)H^{(k)}U_{L}^{\dagger}\left(\theta\right)-\mathrm{diag}\left[U_{L}\left(\theta\right)H^{(k)}U_{L}^{\dagger}\left(\theta\right)\right]\right\|_{F}$ over the ansatz $U_{L}(\theta)$ at each step. This optimization is approximately achieved using the gradient descent method. The numerical simulations are mainly performed using MindSpore Quantum [44].
Ground energy estimation–Estimating expectation values for Hamiltonians is one of the most crucial tasks in state learning, and numerous algorithms have been developed for this purpose  [39, 13, 18, 20]. Sparse Hamiltonians play important roles in quantum computation, chemistry, and simulation [45, 46, 47]. Here, we demonstrate the performance of our algorithms by estimating the ground state energy of a randomly generated sparse Hamiltonian and comparing the results with some existing Pauli decomposition methods [13, 18, 20].
The numerical results are shown in Fig. 2. Here, we randomly generate an 8-qubit sparse Hamiltonian $H$ with $64=8^{2}$ non-zero matrix elements and the ground energy of $-1.532$. In Fig. 2(a), we plot the $2$-norm distance between $\widehat{H}$ and $H$ for various numbers of decomposition terms $K$ in Algorithm 2 and different depth $L$ of the quantum ansatz $U_{L}(\theta)$. Note that, using a shallow circuit, we need only $40$ decomposition terms to reduce the norm to at least half of its original value, which is much smaller than the total number of Pauli terms $4^{8}$. We observe that there exist some platforms in these curves, which arise because the observable decomposition algorithm is based on the Frobenius norm, and we display both norms for comparison. We refer to the Appendix G for the curves of $\left\|\widehat{H}-H\right\|_{F}$, which decays exponentially with the number of decomposition terms.
To demonstrate the sample complexity of our algorithms, we first numerically generate the ground state of the sparse Hamiltonian $H$ and then apply our algorithms to estimate its energy. In Fig. 2(b), we plot the error scaling of the estimated ground energies generated by our algorithms with a fixed number of decomposition terms $K=40$ and a fixed ansatz depth $L=5$. We compare our algorithms with the Derandomized algorithm [13], the Overlapped Grouping Measurement (OGM) algorithm [18], and the Composite Locally Biased Classical Shadow (C-LBCS) algorithm [20]. This figure illustrates the advantage of our algorithm in estimating the expectation value of sparse Hamiltonians.
Figure 2: Performance of energy estimation with a sparse Hamiltonian. (a) The distance between the sparse Hamiltonian $H$ and its estimation $\widehat{H}$ as a function of the number of decomposition terms $K$, with varying depth of the quantum ansatz, $L$, represented in different colors. (b) The estimation errors of the ground energy of the sparse Hamiltonian $H$ using our algorithms with $L=5$ and $K=40$, the Derandomized algorithm [13], the OGM algorithm [18], and the C-LBCS algorithm [20]. The error is calculated from $50$ repeated experiments.
Inner product estimation–Our algorithms can be easily generalized to estimate non-Hermitian operators by allowing $\{\Lambda_{k}\}_{k}$ to be complex-valued diagonal matrices. Hereby, we apply our algorithm to estimate the expectation value of a non-Hermitian operator $O=\ket{1}\ket{\phi_{\tau}}\bra{0^{n+1}}$ |. This operator enables the calculation of the inner product between a Slater determinant and a pure quantum state, where |$\ket{\phi_{\tau}}$ represents a $\tau$-Slater determinant. This has applications in algorithms such as the QC-AFQMC algorithm [48]. Here, $\tau$-Slater determinant [49] is defined as |$\ket{\phi_{\tau}}=\tilde{b}_{1}^{\dagger}\tilde{b}_{2}^{\dagger}\ldots\tilde{b}_{\tau}^{\dagger}\ket{0^{n}}$, where $\tilde{b}_{k}=\sum_{j=1}^{n}U_{kj}b_{j}$, $U$ is an $n\times n$ unitary operator, and $\{b_{j}\}_{j}$ are annihilation operators.
[…]
Figure 3: Performance in estimating the inner product of a Slater determinant and a quantum pure state. (a) The distance between the operator $O=\ket{1}\ket{\phi_{1}}\bra{0^{n+1}}$ |, where |$\ket{\phi_{1}}$ is the 1-Slater determinant, and its approximation $\widehat{O}$ as a function of the number of decomposition terms $K$, with varying the depth of the quantum ansatz, $L$, represented in different colors. (b) The estimated errors for the inner product $\braket{\psi}{\phi_{1}}$, where |$\ket{\psi}$ is a random pure quantum state, using our algorithms with $L=5$ and $K=40$, the Derandomized algorithm [13], the OGM algorithm [18], and the C-LBCS algorithm [20]. The error is calculated from $50$ repeated experiments.
V Lower bound for sample complexity
This work aims to harness the power of shallow circuits for observable estimation. The numerical results presented in the previous section show that performance is closely tied to circuit depth. This naturally raises the question of whether this dependence stems from our algorithms or if there are inherent limitations to the performance of shallow circuits. In this section, we aim to conduct a protocol-independent analysis of shallow circuit performance. Such an analysis will not only help us evaluate our algorithms but also deepen our understanding of the capabilities of shallow quantum circuits in state learning tasks.
Here, we establish a fundamental lower bound on the number of copies of the quantum state $\rho$ required to accurately estimate the value of $\text{Tr}\left(\rho H\right)$. This lower bound holds for all single-copy adaptive strategies that use the given parameterized circuit $U_{L}(\theta)$ without ancillary systems. Our approach builds on the techniques introduced by Chen et al. [50]. While, their work does not cover scenarios where a limited quantum ansatz is utilized.
Theorem 2 (Sample Complexity Lower Bound).
Suppose $H\in\mathbb{C}^{2^{n}\times 2^{n}}$ is a Hermitian operator we aim to estimate. Denote $\mathcal{S}:=\{U_{L}(\theta)^{\dagger}\ket{b}\mid\forall\theta,b\in\{0,1\}^{n}\}$. Then for any single-copy adaptive strategy relying solely on the parameterized circuit $U_{L}(\theta)$ executed on $\rho$, the sample complexity $T$ for measuring $\text{Tr}\left(\rho H\right)$ to accuracy $\epsilon$, for arbitrary unknown state $\rho$, with success probability of at least $2/3$, is lower bounded by
$T=\Omega\left(\frac{\text{Tr}\left(H_{0}^{2}\right)^{2}}{\epsilon^{2}\delta(H_{0})4^{n}}\right),$ (2)
where $H_{0}=H-\frac{\text{Tr}\left(H\right)}{2^{n}}\mathbb{I}_{2^{n}}$ denotes the traceless part of $H$ and $\delta(H_{0}):=\underset{\ket{\psi}\in\mathcal{S}}{\sup}\bra{\psi}H_{0}\ket{\psi}^{2}$.
Here, the dependence of this lower bound on the capability of the circuit is mainly reflected on $\delta(H_{0})$, which increases with the capability of the quantum devices. When the quantum device can prepare all the eigenstates of $H$, $\delta(H_{0})$ reaches its maximum value. For example, when $H=U_{L}(\theta)^{\dagger}PU_{L}(\theta)$, where $P$ is a Pauli operator, we have $\delta(H_{0})=1$. In this case, our lower bound becomes $\Omega\left(\frac{1}{\epsilon^{2}}\right)$, given that $\text{Tr}\left(H_{0}^{2}\right)^{2}=4^{n}$. This lower bound is tight in this case because our parameterized circuit can directly diagonalize $H$. We provide the proof in the Appendix E.
VI Conclusion and discussion
We propose a framework for estimating arbitrary linear properties of a quantum state using available quantum circuits. Our framework takes into account both the limitations of quantum devices and the information of the estimated properties, covering some existing Pauli decomposition protocols [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. Based on this framework, we propose a greedy algorithm that uses a classical optimization algorithm to find an optimized decomposition of the observable $H$ with $L$-depth quantum parameterized circuits such that $H\approx\sum_{k=1}^{K}U_{L}(\theta_{k})^{\dagger}\Lambda_{k}U_{L}(\theta_{k})$. We then sample the decomposition terms using the importance sampling technique. Consequently, the expectation value of $\text{Tr}\left(\rho H\right)$ can be estimated by performing an $L$-depth quantum parameterized circuit sampled from the set $\{U_{L}(\theta_{k})\}_{k=1}^{K}$, followed by collecting the corresponding diagonal matrix $\Lambda_{k}$ with computational basis measurements. Our algorithms can be easily generalized to other circuit structures and non-Hermitian operators. We numerically verify our algorithm by applying it to calculate the ground energy of a given sparse Hermitian and the inner product of a Slater determinant and a pure state.
The performance of our algorithm heavily depends on the decomposition of the target observable. While our greedy decomposition algorithm has been numerically demonstrated to be efficient compared to some conventional methods, it is not guaranteed to be optimal. Therefore, a key future direction is to develop more efficient classical algorithms for optimizing the decomposition process, such as those based on machine learning. Additionally, our current sample complexity lower bound is loose for low-rank observables, such as fidelity. To address this, we could modify the state discrimination tasks, which form the basis of the proof, to establish a tighter bound. It is also worth exploring the extension of our framework to estimate nonlinear properties and incorporate quantum memories.
In addition to minimizing the sample complexity, we aim to reduce the number of required hardware configurations to estimate a target observable. In practical scenarios, repeating experiments with a fixed configuration is relatively straightforward across various platforms, while changing the configuration often requires additional resources. This goal is equivalent to minimizing the number $K$ in the decomposition $H=\sum_{k=1}^{K}U_{k}^{\dagger}\Lambda_{k}U_{k}$, where the $\{U_{k}\}_{k}$ are realizable by the hardware. Thus, an important future direction is to explore the relationship between the number of decomposition terms $K$ and the capability of quantum circuits.
VII Acknowledgement
We would like to acknowledge Zijian Zhang for the valuable discussions regarding the code from their work. We would like to thank Xiao Yuan, Zhenyu Du, Jens Eisert, and Hong-Ye Hu for the helpful discussions. We acknowledge the MindSpore Quantum framework as the platform for our numerics. ZL and XY acknowledge the support from the National Natural Science Foundation of China Grant No. 12174216 and the Innovation Program for Quantum Science and Technology Grant No. 2021ZD0300804 and No. 2021ZD0300702. FW is supported by BMSTC and ACZSP Grant No. Z221100002722017.
Appendix A Proof of Theorem 1
Proof of Theorem 1.
Suppose there exists a set $\{\theta^{(k)},\Lambda_{k}\}_{k=1}^{K}$ such that $\left\|H-\widehat{H}\right\|_{2}\leq\epsilon_{1}$ with $\widehat{H}=\sum_{k=1}^{K}U_{L}(\theta^{(k)})^{\dagger}\Lambda_{k}U_{L}(\theta^{(k)})$, then
$\left|\text{Tr}\left(\rho\widehat{H}\right)-\text{Tr}\left(\rho H\right)\right|\leq\left\|\widehat{H}-H\right\|_{2}\leq\epsilon_{1},$ (3)
where the first inequality holds since $\text{Tr}\left(\rho A\right)\leq\left\|A\right\|_{2}$ for any matrix $A$.
It remains to show that the estimation $v$ generated by Algorithm 1 in combination with Algorithm 2 and important sampling is an unbiased estimation of $\text{Tr}\left(\rho\widehat{H}\right)$ and satisfies
$\Pr\left[{\left|v-\text{Tr}\left(\rho\widehat{H}\right)\right|\leq\epsilon_{2}}\right]\geq 1-\delta$ (4)
where $\delta$ represents the failure probability. Combining inequalities (3) and (4), we obtain
$\left|v-\text{Tr}\left(\rho H\right)\right|\leq\left|v-\text{Tr}\left(\rho\widehat{H}\right)\right|+\left|\text{Tr}\left(\rho\widehat{H}\right)-\text{Tr}\left(\rho H\right)\right|\leq\epsilon_{1}+\epsilon_{2},$ (5)
with a success probability of $1-\delta$. Let $v_{t}=\frac{\Lambda_{k}(b_{t})}{p_{k}}$ be an estimation generated from a single sample, as shown in Step (1) of Algorithm 1. The expectation of $v_{t}$ can then be expressed as
$\mathbb{E}\left[v_{t}\right]$ $=\sum_{k=1}^{K}\frac{1}{p_{k}}p_{k}\mathbb{E}[\Lambda_{k}(b_{t})]$ (6)
$=\sum_{k=1}^{K}\text{Tr}\left(U_{L}(\theta^{(k)})\rho U_{L}(\theta^{(k)})^{\dagger}\Lambda_{k}\right)$ (7)
$=\text{Tr}\left(\rho\widehat{H}\right).$ (8)
Then $\mathbb{E}\left[v\right]=\mathbb{E}\left[v_{t}\right]=\text{Tr}\left(\rho\widehat{H}\right)$. Next, we will bound the variance of $v_{t}$ by noting that
$\mathrm{Var}(v_{t})$ $\leq\mathbb{E}\left[v_{t}^{2}\right]$ (9)
$=\sum_{k=1}^{K}\frac{p_{k}}{p_{k}^{2}}\text{Tr}\left(\rho U_{L}(\theta^{(k)})^{\dagger}\Lambda_{k}^{2}U_{L}(\theta^{(k)})\right)$ (10)
$\leq\sum_{k=1}^{K}\frac{1}{p_{k}}\left\|\Lambda_{k}\right\|_{2}^{2}$ (11)
$=\sum_{k=1}^{K}\frac{\sum_{j=1}^{K}\left\|\Lambda_{j}\right\|_{2}}{\left\|\Lambda_{k}\right\|_{2}}\left\|\Lambda_{k}\right\|_{2}^{2}$ (12)
$=\left\|\bm{\Lambda}\right\|_{1}^{2}$ (13)
where $\bm{\Lambda}$ is a $K$-dimensional vector with the $k$-th entry being $\bm{\Lambda}(k)=\left\|\Lambda_{k}\right\|_{2}$. According to Lemma 1, with $T=\mathcal{O}\left(\frac{\left\|\bm{\Lambda}\right\|_{1}^{2}\log(1/\delta)}{\epsilon_{2}^{2}}\right)$ samplings, the estimation error for $v$ can be be bounded to $\epsilon_{1}+\epsilon_{2}$ in total with a success probability $1-\delta$.
Lemma 1 (Ref. [51]).
Suppose we aim to estimate the expectation value $\mathbb{E}[R]$ of a random variable $R$ with finite variance, using independently drawn samples $R_{1},\cdots,R_{M}$ from its distribution. Then, by employing the median of means estimation method and taking $M=\mathcal{O}\left(\log\left(\delta^{-1}\right)\operatorname{Var}[R]\epsilon^{-2}\right)$, we can guarantee
$\Big{|}\operatorname{\mathbf{MedianOfMeans}}(R_{1},\cdots,R_{M})-\mathbb{E}[R]\Big{|}\leq\epsilon$ (14)
with a success probability of at least $1-\delta$.
Below, we will prove that Algorithm 2 can be conclusively terminated within a finite number of steps, denoted as $K$. Considering one step of parameterized searching, the goal is to minimize the distance $\left\|H-U_{L}\left(\theta\right)^{\dagger}\Lambda U_{L}\left(\theta\right)\right\|_{F}=\left\|U_{L}\left(\theta\right)HU_{L}\left(\theta\right)^{\dagger}-\Lambda\right\|_{F}$. Since $\Lambda$ is a real diagonal matrix, the optimal choice is $\Lambda=\mathrm{diag}\left[U_{L}\left(\theta\right)HU_{L}\left(\theta\right)^{\dagger}\right]$. This choice is optimal because the Frobenius norm quantifies the sum of the squares of all the matrix elements, and choosing $\Lambda$ in this way eliminates all the diagonal terms of $U_{L}\left(\theta\right)HU_{L}\left(\theta\right)^{\dagger}$. Furthermore, with this choice of $\Lambda$, one can prove that $\left\|H-U_{L}\left(\theta\right)^{\dagger}\Lambda U_{L}\left(\theta\right)\right\|_{F}=\left\|U_{L}\left(\theta\right)HU_{L}\left(\theta\right)^{\dagger}-\Lambda\right\|_{F}\leq\left\|U_{L}\left(\theta\right)HU_{L}\left(\theta\right)^{\dagger}\right\|_{F}=\left\|H\right\|_{F}$. This means that when the optimization algorithm functions ideally, a better approximation of $H$ can always be achieved by increasing the iteration steps.
Furthermore, we can prove that it is always possible to find a $\theta$ such that $\left\|H-U_{L}\left(\theta\right)^{\dagger}\Lambda U_{L}\left(\theta\right)\right\|_{F}<\left\|H\right\|_{F}$, provided our parameterized circuit is tomography-complete. If this is not the case, then $\mathrm{diag}[U_{L}\left(\theta\right)HU_{L}\left(\theta\right)^{\dagger}]=0$ must hold for all choices of $\theta$. If $H\neq 0$, this implies the existence of a non-zero observable that cannot be estimated using our parameterized circuits, which contradicts our assumption. We can further prove that if $\theta$ is the optimal value we can find, then $\left\|H\right\|_{F}-\left\|H-U_{L}\left(\theta\right)^{\dagger}\Lambda U_{L}\left(\theta\right)\right\|_{F}\geq\frac{1}{4^{N}}\left\|H\right\|_{F}$. This is because we assume that $U_{L=0}(\theta)$ can realize the tensor product of arbitrary single-qubit gates, thereby enabling the estimation of arbitrary Pauli observables. Given the Pauli decomposition of the observable $H=\sum_{i}\alpha_{i}P_{i}$, one can at least choose $U_{L}\left(\theta\right)^{\dagger}\Lambda U_{L}\left(\theta\right)=\alpha_{j}P_{j}$, where $\alpha_{j}$ is the largest coefficient. In this case, $\left\|H\right\|_{F}-\left\|H-\alpha_{j}P_{j}\right\|_{F}\geq\frac{1}{4^{N}}\left\|H\right\|_{F}$. Considering that the greedy algorithm can achieve better performance than simply picking the largest Pauli term, we can demonstrate that $\left\|H\right\|_{F}-\left\|H-U_{L}\left(\theta\right)^{\dagger}\Lambda U_{L}\left(\theta\right)\right\|_{F}\geq\frac{1}{4^{N}}\left\|H\right\|_{F}>0$. Thus, our algorithm can terminate in a finite number of steps. ∎
Appendix B Qubit-commuting decomposition
We use the symbol $Q\triangleright P$ to indicate that $P$ covers $Q$, meaning that each $i$-th qubit of $Q$, denoted as $Q_{i}$, is either equal to the identity or to $P_{i}$. Let the operator $H=\sum_{i=1}^{m}\alpha_{i}Q_{i}$, where $\{Q_{i}\}_{i}$ are Pauli operators and $\{\alpha_{i}\}_{i}$ are real coefficients. Let $e_{1},\ldots,e_{K}$ be (overlapped) qubit-commuting sets such that $e_{i}\subseteq\{Q_{1},\ldots,Q_{m}\}$. We claim that there exists a Pauli operator $P^{(i)}$ to cover all elements of the qubit-commuting set $e_{i}$ for $i\in\{1,\ldots,K\}$.
The eigendecomposition for $P^{(i)}$ can be represented as $P^{(i)}=U_{i}^{\dagger}O_{i}U_{i}$ where $U_{i}$ is a unitary and $O_{i}$ is diagonal. For $n$-qubit tensor product unitary $A_{j}$, denote the $l$-th qubit representation as $A_{jl}$, i.e., $A_{j}=\otimes_{l=1}^{n}A_{jl}$. Then we have
$H$ $=\sum_{i=1}^{m}\alpha_{i}Q_{i}$ (15)
$=\sum_{i=1}^{K}\sum_{Q_{j}\in e_{i}}\alpha_{j}Q_{j}$ (16)
$=\sum_{i=1}^{K}\sum_{Q_{j}\in e_{i}}\alpha_{j}\bigotimes_{l=1}^{n}Q_{jl}$ (17)
$=\sum_{i=1}^{K}\sum_{Q_{j}\in e_{i}}\alpha_{j}U_{i}^{\dagger}\left(\bigotimes_{l=1}^{n}D_{jl}\right)U_{i}$ (18)
$=\sum_{i=1}^{K}U_{i}^{\dagger}\left(\sum_{Q_{j}\in e_{i}}\alpha_{j}D_{j}\right)U_{i}$ (19)
$=\sum_{i=1}^{K}U_{i}^{\dagger}\Lambda_{i}U_{i}$ (20)
where $D_{jl}$ equals the single-qubit identity $\mathbb{I}_{2}$ if $Q_{jl}=\mathbb{I}_{2}$ and $D_{jl}=O_{jl}$ otherwise, Eq. (18) holds since $Q_{j}\triangleright P^{(i)}$ for any $Q_{j}\in e_{i}$, and Eq. (20) holds by setting $\Lambda_{i}=\sum_{Q_{j}\in e_{i}}\alpha_{j}D_{j}$.
Appendix C Review of Qubit-Commuting Random Measurements
Let the Pauli operator expansion of a Hermitian operator $H=\sum_{i=1}^{m}\alpha_{i}Q_{i}$. We group all $m$ Pauli operators $\left\{Q_{i}\right\}$ into overlapped sets $e_{1},e_{2},\ldots,e_{K}$, where all Pauli operators in each $e_{i}$ are commute to each other. There exists a Pauli operator $P^{(i)}$ such that all Pauli operators $Q_{j}$ in $e_{i}$ satisfy $Q_{j}\triangleright P^{(i)}$ for any $i\in[K]$. For instance, when $e=\left\{Q_{1},Q_{2}\right\}$, $Q_{1}=Z_{1}X_{2}I_{3}$ and $Q_{2}=Z_{1}I_{2}Y_{3}$, we have $P=Z_{1}X_{2}Y_{3}$.
In the OGM algorithm [18] and other grouping measurement algorithms [52, 53], sets $e_{1},\ldots,e_{K}$ and measurements $P^{(1)},\ldots,P^{(K)}$ are explicitly generated with associated probabilities $p_{1},\ldots,p_{K}$. In other random measurement algorithms, while the groups are not explicitly generated, the analysis in Ref. [18] shows that they implicitly use the same idea. Some algorithms, such as the Derandomized algorithm proposed by Huang et al. [13], generate the counts of $P^{(1)},\ldots,P^{(K)}$ when the total number of measurements $T$ is specified, essentially creating a derandomized version of the probabilistic sampling.
Since the Clifford group stabilizes the Pauli group, any Pauli operator can be decomposed into $P=U^{\dagger}\Lambda U$, where $U$ is the tensor product of local Clifford gates. Measuring a quantum state $\rho$ with a Pauli operator $P$ is then equivalent to performing a local Clifford gate $U$, followed by measurement in the computational basis to collect the associated eigenvalues of $P$ [39]. Consequently, qubit-commuting random measurement schemes involve only local Clifford circuits.
Appendix D Review of Commuting Measurements
In this section, we discuss how to decompose $H=\sum_{j=1}^{m}\alpha_{j}Q_{j}$ into $\sum_{k=1}^{M}U_{k}^{\dagger}\Lambda_{k}U_{k}$ where $Q_{j}$ are Pauli operators, $U_{k}$ are Clifford circuits and $M$ is the number of commuting groups of $\left\{Q_{j}\right\}_{j=1}^{m}$ [11].
Let $H=\sum_{k=1}^{K}H_{k}$ where $H_{k}=\sum_{Q_{j}\in g_{k}}\alpha_{j}Q_{j}$ where any two Pauli operators $P_{i},P_{j}$ in set $g_{k}$ are commute to each other. To generate the $U_{k}$ and $\Lambda_{k}$ for $H_{k}$, we start by finding sets $\mathcal{G}=\left\{\vartheta_{1},\ldots,\vartheta_{n}\right\}$ and $\mathcal{Q}=\left\{\sigma_{1},\ldots,\sigma_{n}\right\}$ where $\vartheta_{i}$ are Pauli operators and $\sigma_{j}$ are single-qubit Pauli operators such that
  • $[\vartheta_{i},\vartheta_{j}]=0$ for any $i,j\in[n]$;
  • $[\vartheta_{i},Q_{l}]=0$ for any $i\in[n],Q_{l}\in g_{k}$;
  • $\left\{\vartheta_{i},\sigma_{j}\right\}=0$ for any $i\neq j\in[n]$;
  • $[\vartheta_{i},\sigma_{i}]=0$ for any $i\in[n]$;
  • $[\sigma_{i},\sigma_{j}]=0$ for any $i,j\in[n]$.
According to Ref. [54], we can give the set $\mathcal{G}$, which serves as the generator set for Pauli operators in $g_{k}$, meaning that all $Q_{j}$ can be expressed as a product of some elements $\vartheta_{i}$ in $\mathcal{G}$. It can also be easily proved that such set $\mathcal{Q}$ also exists. Let $U_{k}^{\prime}=\prod_{l=1}^{n}\frac{1}{\sqrt{2}}\left(g_{l}+\sigma_{l}\right)$. Suppose $\sigma_{l}=C_{l}^{\dagger}D_{l}C_{l}$. Let $L_{k}=\prod_{l=1}^{n}C_{l}$ and $U_{k}=L_{k}U_{k}^{\prime}$. Then we have $U_{k}^{\prime}g_{l}U_{k}^{\prime\dagger}=\sigma_{l}$, hence $U_{k}g_{l}U_{k}^{\dagger}=L_{k}\sigma_{l}L_{k}^{\dagger}=D_{l}$, and
$U_{k}H_{k}U_{k}^{\dagger}=\sum_{Q_{j}\in g_{k}}\alpha_{j}\prod_{\vartheta_{l}\in S_{j}}D_{l},$
where $Q_{j}=\prod_{\vartheta_{l}\in S_{j}}g_{l}$. Such $S_{j}$ exists by the definition of $\left\{\vartheta_{i}\right\}_{i}$. Let $\Lambda_{k}=\sum_{Q_{j}\in g_{k}}\alpha_{j}\prod_{\vartheta_{l}\in S}D_{l}$, then $H=\sum_{k=1}^{K}U_{k}^{\dagger}\Lambda_{k}U_{k}$.
Appendix E Proof of Theorem 2
We utilize the methods outlined in Ref. [50] to demonstrate our theorem. WLOG, we can assume $\text{Tr}\left(H\right)=0$. Denote $H=H_{1}=-H_{2}$. Consider the following two state sets:
  • Set 1: $\{\mathbb{I}/2^{n}\}$,
  • Set 2: $\{\left(\mathbb{I}+3aH_{1}\right)/2^{n},\left(\mathbb{I}+3aH_{2}\right)/2^{n}\}$, with equal probability,
where we take $a\leq\frac{1}{3\left\|H\right\|_{2}}$ to make sure matrices in Set 2 are valid density matrices, where $\left\|H\right\|_{2}$ is two-norm (spectral norm) of $H$.
Consider the state discrimination task associated with these two state sets: a referee tosses a fair coin to select a set, and then chooses a state $\rho$ from the selected set according to the distribution. Then the referee gives $T$ copies of $\rho$ to us and our task is to decide which set $\rho$ is from, with $T$ copies of $\rho$. What we have is a $n$-qubit quantum device of limited capability, whose all possible output states when acting on |$\ket{0^{n}}$ form a set $\mathcal{S}$.
Note that $\text{Tr}\left(H\frac{\mathbb{I}}{2^{n}}\right)=0$, $\text{Tr}\left(H\frac{\mathbb{I}+3aH_{1}}{2^{n}}\right)=\frac{3a\text{Tr}\left(H^{2}\right)}{2^{n}}$ and $\text{Tr}\left(H\frac{\mathbb{I}+3aH_{2}}{2^{n}}\right)=-\frac{3a\text{Tr}\left(H^{2}\right)}{2^{n}}$. One possible way to accomplish the state discrimination task is to measure $H$ to precision $\epsilon=\frac{a\text{Tr}\left(H^{2}\right)}{2^{n}}$ and check whether the approximate expectation value lies in $\big{[}-\frac{a\text{Tr}\left(H^{2}\right)}{2^{n}},\frac{a\text{Tr}\left(H^{2}\right)}{2^{n}}\big{]}$. If it holds, we assign $\rho$ to Set 1, otherwise we assign $\rho$ to Set 2.
Thus if we can use this quantum device to measure $\text{Tr}\left(H\rho\right)$ with an accuracy of $\epsilon=\frac{a\text{Tr}\left(H^{2}\right)}{2^{n}}$ and a success probability of at least $2/3$ for any given $\rho$, then we can accomplish the state discrimination task with the same success probability of at least $2/3$ using this quantum device.
In the following, we show the sample complexity $T$ for accomplishing the state discrimination task by this quantum device with a success probability of at least $2/3$, is lower bounded by $T\geq\Omega\left(\frac{\text{Tr}\left(H^{2}\right)^{2}/4^{n}}{\epsilon^{2}\delta(H)}\right)$, thus this sample complexity for observable measurement is also lower bounded by this quantity.
When the quantum device is capable of performing arbitrary POVM measurement, any state discrimination protocol can be associated with a decision tree shown by Fig. 4 in [55]. Now since our quantum device can only perform parameterized unitary $U_{L}(\theta)$ and computational basis measurement, the experiment in each node of the decision tree is changed to $\{\alpha_{i}(\theta)\}_{i=1}^{2^{n}}$, where $\alpha_{i}(\theta)$ is the $i$-th column of $U_{L}(\theta)^{\dagger}$.
Given any state discrimination protocol using this quantum device, and a chosen state $\rho$ given by the referee, we use $l$ to label the leaves of the tree. Let $p_{\rho}$ be a probability distribution on the leaves, such that $p_{\rho}(l)$ equals the probability of arriving at the leaf $l$. For probability distributions $p=\{p_{1},\cdots,p_{N}\}$ and $q=\{q_{1},\cdots,q_{N}\}$, we define the total variation between them by $\operatorname{TV}(p,q)=\frac{1}{2}\sum_{i=1}^{N}\left|p_{i}-q_{i}\right|$. To accomplish the state discrimination task with a success probability of at least $2/3$, we need to make the depth $T$ of the decision tree, i.e., the sample complexity big enough to ensure [50]
$\operatorname{TV}\left(\mathbb{E}_{x}p_{\rho_{x}},p_{\mathbb{I}/2^{n}}\right)\geq\frac{1}{3}.$ (21)
For a leaf $l$ in the decision tree, there is a unique path from the root node $n_{0}$ to the leaf node $n_{T}=l$. Denote the nodes on the path by $n_{0},\cdots,n_{T}$ and the edges between them by $e_{1},\cdots,e_{T}$. For $t=1,\cdots,T$, denote the projector associated with $e_{t}$ by $|\alpha_{e_{t}}(\theta_{n_{t-1}})\rangle\langle\alpha_{e_{t}}(\theta_{n_{t-1}})|$. To calculate the total variation in Eq. (21), we calculate the following quantity:
$\frac{\mathbb{E}_{x}p_{\rho_{x}}(l)}{p_{\mathbb{I}/2^{n}}(l)}$ $=\mathbb{E}_{x}\prod_{t=1}^{T}\left(\frac{1/2^{n}+1/2^{n}3a\bra{\alpha_{e_{t}}(\theta_{n_{t-1}})}H_{x}\ket{\alpha_{e_{t}}(\theta_{n_{t-1}})}}{1/2^{n}}\right)$ (22)
$=\mathbb{E}_{x}\prod_{t=1}^{T}\Big{(}1+3a\bra{\alpha_{e_{t}}(\theta_{n_{t-1}})}H_{x}\ket{\alpha_{e_{t}}(\theta_{n_{t-1}})}\Big{)}$
$=\mathbb{E}_{x}\exp\left[\sum_{t=1}^{T}\log\Big{(}1+3a\bra{\alpha_{e_{t}}(\theta_{n_{t-1}})}H_{x}\ket{\alpha_{e_{t}}(\theta_{n_{t-1}})}\Big{)}\right]$
$\geq\exp\left[\sum_{t=1}^{T}\mathbb{E}_{x}\log\Big{(}1+3a\bra{\alpha_{e_{t}}(\theta_{n_{t-1}})}H_{x}\ket{\alpha_{e_{t}}(\theta_{n_{t-1}})}\Big{)}\right]$
$=\exp\left[\sum_{t=1}^{T}\frac{1}{2}\log\Big{(}1-9a^{2}\bra{\alpha_{e_{t}}(\theta_{n_{t-1}})}H\ket{\alpha_{e_{t}}(\theta_{n_{t-1}})}^{2}\Big{)}\right]$
$\geq\exp\left[-9a^{2}\sum_{t=1}^{T}\bra{\alpha_{e_{t}}(\theta_{n_{t-1}})}H\ket{\alpha_{e_{t}}(\theta_{n_{t-1}})}^{2}\right]$
$\geq\exp\left[-9a^{2}T\delta(H)\right]$
$\geq 1-9Ta^{2}\delta(H)$
where the second inequality holds because $\log(1-x)\geq-2x$ for all $x\in[0,0.79]$. We can then use it to estimate the total variation as
[…]
To achieve a success probability of at least $2/3$, we need to ensure
[…]
thus we have
$T\geq\Omega\left(\frac{1}{a^{2}\delta(H)}\right)=\Omega\left(\frac{\text{Tr}\left(H^{2}\right)^{2}/4^{n}}{\epsilon^{2}\delta(H)}\right).$ (25)
Appendix F Preliminary to inner product calculation
[…]
Since the observable $H$ is not Hermitian and $\text{Tr}\left(\rho H\right)$ cannot be directly approximated with existing random measurement algorithms. When executing the existing algorithm to calculate the inner product, we decompose the observable $H$ into two Hermitians $H_{1}:=(H+H^{\dagger})/2$ and $H_{2}:=-\mathrm{i}(H-H^{\dagger})/2$, then $H=H_{1}+\mathrm{i}H_{2}$, and it can then be approximated in two parts. The total error $\varepsilon=\left|\hat{v}-\text{Tr}\left(\rho H\right)\right|=\sqrt{\left(\hat{v}_{1}-\text{Tr}\left(\rho H_{1}\right)\right)^{2}+\left(\hat{v}_{2}-\text{Tr}\left(\rho H_{2}\right)\right)^{2}}$ where $\hat{v}_{j}$ is the estimation of $\text{Tr}\left(\rho H_{j}\right)$ for $j\in\left\{1,2\right\}$.
Appendix G Supplementary numerics
Here, we present the changes in $\left\|H^{(k)}-H\right\|_{F}$ as $k$ increases. As illustrated in Fig. 4, the Frobenius distance between $H^{(k)}$ and $H$ decays exponentially.
Figure 4: The variations in Frobenius distance as the number of decomposition terms increases for (a) a sparse Hermitian $H$ and (b) a non-Hermitian operator $O$ associated with a Slater determinant.